# ToDRE: Visual Token Pruning via Diversity and Task Awareness for Efficient Large Vision-Language Models

### v1

@@ 4 ToDRE: Two-stage Visual Token Pruning via Diversity and Task Awareness
Figure 4: Overall framework of ToDRE. Diversity-driven Token Selection first selects a pivot token from global thumbnail or video frames using [CLS]-based attention, then applies a greedy $k$-center expansion strategy to retain a diverse set of $k$ visual tokens. Relevance-driven Visual Token Reduction dynamically identifies a layer at which all visual tokens are pruned. A layer is selected for pruning if both the visual-to-text and text-to-visual attention ratios fall below a predefined threshold. $E^{G}_{v}$, $E^{C}_{v}$, and $E^{F}_{v}$ respectively denote the embeddings of thumbnail, local crops, and video frames.
@@
Building on the preliminary analysis (Section 3), we introduce ToDRE, a two-stage, training-free, and plug-and-play visual token compression framework (see Figure 4). ToDRE utilizes a similarity-guided greedy search in the LLM’s embedding space to select a maximally diverse subset of visual tokens, followed by an adaptive task-relevance-based pruning mechanism within the LLM’s decoder. Next, we elaborate each stage in detail.
4.1 Stage 1: Diversity-driven Visual Token Selection
@@
To obtain a maximally diverse subset of visual tokens, we adopt a greedy k-center algorithm [16] consisting of two steps: (1) initializing a retention set by selecting the initial pivot token, and (2) iteratively adding the token whose maximum similarity to the current set is smallest. Full pseudocode of our proposed token retention algorithm is provided in Algorithm 1.
@@
To determine the initial pivot, we leverage the [CLS] attention from the last layer of the vision encoder [45] as an importance indicator. The attention from the [CLS] token $\bm{z}_{[\mathrm{CLS}]}\in\mathbb{R}^{d}$ to other visual tokens $\bm{Z}_{v}\in\mathbb{R}^{n\times d}$ is calculated as:
$\bm{q}_{[\mathrm{CLS}]}=\bm{z}_{[\mathrm{CLS}]}\bm{W}_{Q},\quad\bm{K}_{v}=\bm{Z}_{v}\bm{W}_{K},\quad\bm{a}_{[\mathrm{CLS}]}=\operatorname{Softmax}\left(\frac{\bm{q}_{[\mathrm{CLS}]}\bm{K}_{v}^{T}}{\sqrt{d}}\right),$ (3)
where $n$ is the length of the visual token sequence; $d$ is the hidden state size of vision encoder; $\bm{W}_{Q}\in\mathbb{R}^{d\times d}$ and $\bm{W}_{K}\in\mathbb{R}^{d\times d}$ represent the weight matrices for queries and values, respectively.
As shown in Figure 4-(a), pivot token selection proceeds as follows: (1) Image Inputs with AnyRes [36] Support: In this case, LVLM yields one global thumbnail $G$ along with several local crops $C$. We compute the [CLS] attention score for each token in the global thumbnail and choose the token with the highest score as the pivot, since it captures the most comprehensive global information. (2) Image Inputs without AnyRes Support: The pivot token is selected from all visual tokens of the original image, using the same [CLS]-based criterion. (3) Video Inputs: We first identify, for each frame, the visual token with the highest [CLS] attention. The final pivot token is then selected as the one with the highest score among these frame-wise candidates.
Greedy $k$-center Expansion.
The expansion process begins with the designated pivot token. In each iteration, a new token index $c^{(t)}$ is selected in a greedy manner, which can be formulated as follows:
$c^{(t)}=\operatorname*{arg\,min}_{v\in V\setminus\mathcal{C}^{(t-1)}}\left[\max_{c\in\mathcal{C}^{(t-1)}}(s(\mathbf{x_{v}},\mathbf{x_{c}}))\right],$ (4)
where $\mathbf{x_{v}}$ and $\mathbf{x_{c}}$ represent visual tokens of index $v$ and $c$, $\mathcal{C}^{({t-1})}$ denotes the retention set from the previous iteration, and $V\setminus\mathcal{C}^{(t-1)}$ is the remaining candidate tokens. The similarity $\mathrm{s}(\mathbf{x_{v}},\mathbf{x_{c}})$ between two tokens is measured with cosine similarity:
$\mathrm{s}(\mathbf{x_{v}},\mathbf{x_{c}})=\frac{\mathbf{x_{v}}\cdot\mathbf{x_{c}}}{\|\mathbf{x_{v}}\|\|\mathbf{x_{c}}\|}.$ (5)
The algorithm selects the token that has the smallest maximum similarity to the current set to ensure maximal dissimilarity. The selected token is then added to $\mathcal{C}$, and the similarity scores are updated accordingly. This process is repeated until $k$ diverse tokens (e.g., $k=288$, corresponding to 10% of visual tokens) are retained, constructing the final retention set $\mathcal{C}$ as:
@@
Once the retention set is finalized, all remaining candidate visual tokens are discarded. The retained visual tokens, along with all the text tokens, are subsequently fed into the LLM decoder for inference.
4.2 Stage 2: Relevance-driven Visual Token Reduction
While strategies involving partial or multi-stage pruning could be further applied, we argue that such strategies are unnecessary, since the majority of visual tokens have already been removed at Stage 1. In contrast to VTW [35], which relies on post hoc KL-divergence comparisons to determine the optimal pruning layer—a method that is indirect and non-intuitive—we propose a forward-pass metric based on cross-modal attention that directly identifies the most appropriate layer in LLM for token removal based on actual token interaction. As shown in Figure 4-(b), all visual tokens are removed after this selected layer.
Specifically, let $L$ be the number of decoder layers of LLM. Based on our empirical observation (Figure 2) that deeper layers exhibit limited cross-modal interaction, we compute cross-modal attention ratios only at a few selected layers in the later prefilling stages of the model. Since these attention ratios tend to remain stable across consecutive deeper layers, computing them at every layer would introduce unnecessary overhead. In our implementation, we select layers located at fractional depths $L/{2}$, $85L/$ 8, $86L/$ 8 and $87L/$ 8. A more detailed ablation of layer selection can be found in Section C.2. At each selected layer $\ell$, we compute two cross-modal attention ratios based on average attention probabilities across all attention heads and tokens:
$\alpha_{t\to v}^{(\ell)}=\frac{\sum_{i\in T}\sum_{j\in V}A_{ij}^{(\ell)}}{\sum_{i\in T}\sum_{j\in S\cup V\cup T}A_{ij}^{(\ell)}},\quad\alpha_{v\to t}^{(\ell)}=\frac{\sum_{i\in V}\sum_{j\in T}A_{ij}^{(\ell)}}{\sum_{i\in V}\sum_{j\in S\cup V\cup T}A_{ij}^{(\ell)}},$ (7)
where $A_{ij}^{\ell}$ denotes the softmax-normalized attention weight from query token $i$ to key token $j$ at layer $\ell$; $S$, $V$, and $T$ represent the system prompt, visual, and textual tokens, respectively. To further enhance efficiency, all visual tokens are removed at a certain layer $\ell$ if and only if both $\alpha_{t\to v}^{(\ell)}$ and $\alpha_{v\to t}^{(\ell)}$ are lower than a threshold $\tau$. By removing all visual tokens at this point, the model further avoids redundant visual computation in the remaining prefilling and decoding stages, yielding slight improvements in both efficiency and performance.
@@ 5 Experiments
Our experiments were conducted using LLaVA-NeXT-7B/13B models [37] on a total of twelve benchmarks, including eight for image understanding and four for video understanding. More details on benchmarks and comparing methods can be found in Appendix A.
5.1 Main Results
@@
In Table 1, we report ToDRE’s performance on a range of image-understanding benchmarks at different token-retain ratios. First, under the same setup where 75% of visual tokens are pruned in Stage 1—matching competing methods—ToDRE further removes all remaining visual tokens in Stage 2 and still achieves a 97.3% average score, outperforming the second-best method by 1.7%. Second, under more extreme compression (only 10% visual token retained), ToDRE surpasses the second-best approach by 2.9%. Third, ToDRE also achieves top performance on larger models, reaching 93.4% average scores on the 13B variant—demonstrating strong adaptability across model scales. Note that FastV [10] and SparseVLM [62] are excluded from the 13B comparison, as their pruning strategies, originally tailored for the 7B model, lead to substantial performance degradation when directly transferred to 13B. This further underscores the robustness and transferability of ToDRE.
@@
To further assess ToDRE’s generalization ability, we evaluate it on both short- and long-form video understanding benchmarks. As shown in Table 2, ToDRE outperforms the baseline by 5.0% and 3.3% under the same token-retain ratios used for images, and surpasses the second-best method by 0.4% and 6.9%, respectively. Interestingly, ToDRE even surpasses the baseline model in some cases. We attribute this to the reduced interference from redundant visual tokens, which may otherwise suppress task-relevant information during inference. Similarly, SparseVLM is excluded due to transferability issues, and GlobalCom2 [39] is omitted as it is specifically designed for image-only inputs. In contrast, ToDRE demonstrates broad generalization across both modalities and model scales.
Method MME ScienceQA GQA POPE MMBench MMBench-CN VizWiz VQAv2 Average
Upper Bound, 2880 Tokens
LLaVA-NeXT-7B [37] 1519.6 72.0 64.2 87.7 68.5 59.0 60.6 80.1 100.0%
Ratio=25%, Retain up to 720 Tokens
FastV [10] 1477.3 69.8 60.4 83.1 65.6 55.4 57.2 77.2 95.4%
SparseVLM [62] 1446.1 67.5 60.9 71.0 63.8 55.4 58.6 77.2 93.1%
FasterVLM [61] 1454.6 67.1 61.3 87.2 66.0 56.8 58.4 76.4 96.0%
GlobalCom2 [39] 1468.7 68.1 61.4 87.6 64.0 54.4 58.7 76.6 95.6%
ToDRE (Ours) 1483.3 70.2 61.4 87.7 66.2 58.0 58.4 77.2 97.3%
Ratio=10%, Retain up to 288 Tokens
FastV [10] 1282.9 69.3 55.9 71.7 61.6 53.5 56.1 70.2 88.8%
SparseVLM [62] 1332.2 68.6 56.1 63.2 54.5 52.3 56.2 69.9 86.3%
FasterVLM [61] 1359.2 66.5 56.9 83.6 61.6 55.1 55.6 72.3 91.4%
GlobalCom2 [39] 1365.5 68.7 57.1 83.8 61.8 55.0 57.6 71.8 92.2%
ToDRE (Ours) 1413.1 70.0 60.3 85.9 64.0 57.1 57.8 74.5 95.1%
LLaVA-NeXT-13B [37] 1575.2 71.2 65.4 87.5 70.1 66.0 63.6 81.9 100.0%
Ratio=25%, Retain up to 720 Tokens
FasterVLM [61] 1516.1 71.1 62.3 86.1 67.6 62.1 58.1 76.1 95.6%
GlobalCom2 [39] 1531.2 71.4 62.7 86.5 67.9 61.3 58.2 77.2 96.0%
ToDRE (Ours) 1542.1 71.6 62.9 87.4 68.0 62.0 60.8 78.0 97.0%
Ratio=10%, Retain up to 288 Tokens
FasterVLM [61] 1386.2 70.5 58.1 81.6 61.7 53.5 55.9 77.1 89.1%
GlobalCom2 [39] 1399.5 71.0 58.3 82.4 65.0 56.6 55.6 72.8 90.8%
ToDRE (Ours) 1501.8 71.2 60.7 86.1 65.6 54.4 59.0 75.5 93.4%
Table 1: Performance of training-free token compression methods across eight image-language benchmarks. “Average” denotes the mean performance ratio between each token compression method and the vanilla LLaVA-NeXT-7B. We evaluate all methods at retain ratios of 25% and 10%, with the best results highlighted in bold.
LLaVA-NeXT-7B [37] 0% 2880 33.3 35.7 20.1 42.5 100.0%
FastV [10] 25% 720 32.3 31.2 16.5 40.3 90.4%
FasterVLM [61] 36.9 41.0 19.6 40.5 104.6%
ToDRE (Ours) 33.5 41.4 20.3 43.5 105.0%
FastV [10] 10% 288 30.4 30.4 11.2 38.7 80.8%
FasterVLM [61] 34.3 36.1 19.1 36.9 96.4%
ToDRE (Ours) 33.4 40.6 20.0 42.3 103.3%
5.2 Computational Efficiency
As shown in Table 3, we compare FLOPs, peak memory usage, throughput, and performance across various token pruning methods under a fixed token-retain ratio of 10%. First, ToDRE achieves the highest throughput of 2.8 samples/s on POPE [31], accelerating inference by 1.9$\times$ compared to the vanilla LLaVA-NeXT-7B baseline, while matching the lowest memory usage (13.6 GB) alongside FasterVLM [61]. Second, despite its superior efficiency and memory cost, ToDRE maintains the highest average performance (95.1%), outperforming the second-best method by 2.9%. These results confirm that ToDRE achieves great overall balance among speed, memory, and accuracy. We attribute the slight efficiency gains over FasterVLM and GlobalCom2 (FLOPs $\uparrow$0.3%, throughput $\uparrow$0.1 samples/s) to our second-stage deletion of all remaining visual tokens—an approach rarely adopted in prior work. In addition, as discussed in Section 3.1, because most image and video understanding benchmarks only require the model to answer a single word or short phrase (where $L$ is considerably small), our efficiency gains during the LLM decoding stage are inevitably marginal. However, we expect ToDRE to deliver even greater efficiency benefits in tasks involving longer text generation, since it effectively mitigates the computational burden of visual tokens during LVLM inferences.

### v2

@@ 4 Visual Token Pruning with Token Diversity and Task Relevance
Figure 3: Overall framework of ToDRE. Given the visual and textual inputs, the proposed Diversity-driven Token Selection first selects a pivot token from global thumbnail or video frames with [CLS]-based attention and then performs max-sum diversification to retain a diverse set of $k$ visual tokens. The proposed Relevance-driven Token Reduction then dynamically identifies a pivot decoder layer and prunes all its visual tokens—the layer is identified if its visual-to-text and text-to-visual attention ratios both fall below a predefined threshold $\tau$. $E^{G}_{v}$, $E^{C}_{v}$, and $E^{F}_{v}$ denote the embeddings of thumbnail, local crops, and video frames, respectively.
@@
Building on the preliminary analysis, we introduce ToDRE, a two-stage, training-free, and plug-and-play visual token compression framework (see Figure˜3). ToDRE utilizes a similarity-guided greedy search in the LLM embedding space to select a maximally diverse subset of visual tokens, followed by an adaptive task-relevance-based pruning mechanism within the LLM decoder. Next, we elaborate on each stage in detail.
4.1 Diversity-Driven Token Selection
@@
To obtain a maximally diverse subset of visual tokens, we adopt a greedy max-sum diversification algorithm [10.1145/1526709.1526760] consisting of two steps: (1) initializing a retention set by selecting the initial pivot token, and (2) iteratively adding the token that minimizes its cumulative similarity to the current set. Full pseudocode of our proposed token retention algorithm is provided in Appendix.
@@
To determine the initial pivot, we leverage the [CLS] attention from the last layer of the vision encoder [radford2021clip] as an importance indicator. The attention from the [CLS] token $\bm{z}_{[\mathrm{CLS}]}\in\mathbb{R}^{d}$ to other visual tokens $\bm{Z}_{v}\in\mathbb{R}^{n\times d}$ is calculated as:
$\bm{q}_{[\mathrm{CLS}]}$ $=\bm{z}_{[\mathrm{CLS}]}\bm{W}_{Q},\quad\bm{K}_{v}=\bm{Z}_{v}\bm{W}_{K},$ (3)
$\bm{a}_{[\mathrm{CLS}]}$ $=\operatorname{Softmax}\!\left(\frac{\bm{q}_{[\mathrm{CLS}]}\bm{K}_{v}^{\!\top}}{\sqrt{d}}\right),$
where $n$ is the length of the visual token sequence; $d$ is the hidden state size of vision encoder; $\bm{W}_{Q}\in\mathbb{R}^{d\times d}$ and $\bm{W}_{K}\in\mathbb{R}^{d\times d}$ represent the weight matrices for queries and keys, respectively.
As shown in Figure˜3-(a), pivot token selection proceeds as follows: (1) Image Inputs with AnyRes [liu2024llava15] Support: In this case, LVLM yields one global thumbnail $G$ along with several local crops $C$. We compute the [CLS] attention score for each token in the global thumbnail and choose the token with the highest score as the pivot, since it captures the most comprehensive global information. (2) Image Inputs without AnyRes Support: The pivot token is selected from all visual tokens of the original image, using the same [CLS]-based criterion. (3) Video Inputs: We first identify, for each frame, the visual token with the highest [CLS] attention. The final pivot token is then selected as the one with the highest score among these frame-wise candidates.
For MLLMs without a [CLS] token in their encoders, a random selection strategy is also acceptable, as it yields performance that is nearly comparable to the original approach. We provide a detailed comparison of different pivot token selection strategies in Appendix.
Greedy Max-Sum Diversification.
The expansion starts from the designated pivot. At iteration $t$, we pick a new token index $c^{(t)}$ by minimizing its cumulative similarity to the already selected set:
$c^{(t)}\;=\;\operatorname*{arg\,min}_{v\in V\setminus\mathcal{C}^{(t-1)}}\Bigg[\sum_{c\in\mathcal{C}^{(t-1)}}s(\mathbf{x}_{v},\mathbf{x}_{c})\Bigg],$ (4)
where $\mathbf{x}_{v}$ and $\mathbf{x}_{c}$ denote visual token features with indices $v$ and $c$, and $\mathcal{C}^{(t-1)}$ is the selected set from the previous iteration. The similarity between two tokens is measured with cosine similarity
$s(\mathbf{x}_{v},\mathbf{x}_{c})\;=\;\frac{\mathbf{x}_{v}^{\top}\mathbf{x}_{c}}{\lVert\mathbf{x}_{v}\rVert\,\lVert\mathbf{x}_{c}\rVert}.$ (5)
Equivalently, (4) maximizes the sum of distances if $d(\cdot,\cdot)=1-s(\cdot,\cdot)$. After selecting $c^{(t)}$, we update the cumulative similarities by adding its contribution:
$\forall v\in V\setminus\mathcal{C}^{(t)}:\quad S_{v}^{(t)}\;=\;S_{v}^{(t-1)}\;+\;s(\mathbf{x}_{v},\mathbf{x}_{c^{(t)}}),$ (6)
and mask the chosen index. This greedy procedure repeats until $k$ diverse tokens (e.g., k=288k{=}288, about $10\%$ of visual tokens) are retained, yielding
@@
Finally, all remaining visual tokens are discarded; the retained visual tokens together with all text tokens are fed to the LLM decoder for inference.
4.2 Relevance-Driven Token Compression
While strategies involving partial or multi-stage pruning could be further applied, we argue that such strategies are unnecessary, since the majority of visual tokens have already been removed at Stage 1. In contrast to VTW [lin2025vtw], which relies on post hoc KL-divergence comparisons to determine the optimal pruning layer—a method that is indirect and non-intuitive—we propose a forward-pass metric based on cross-modal attention that directly identifies the most appropriate layer in LLM for token removal based on actual token interaction. As shown in Figure˜3-(b), all visual tokens are removed after this selected layer.
Specifically, let $L$ be the number of decoder layers of LLM. Based on our empirical observation (Figure˜2) that deeper layers exhibit limited cross-modal interaction, we compute cross-modal attention ratios only at a few selected layers in the later prefilling stages of the model. Since these attention ratios tend to remain stable across consecutive deeper layers, computing them at every layer would introduce unnecessary overhead. In our implementation, we select layers located at fractional depth $7L/8$. A more detailed ablation of layer selection can be found in Appendix. At each selected layer $\ell$, we compute two cross-modal attention ratios based on average attention probabilities across all attention heads and tokens:
$\alpha_{t\to v}^{(\ell)}$ $=\frac{\sum_{i\in T}\sum_{j\in V}A_{ij}^{(\ell)}}{\sum_{i\in T}\sum_{j\in S\cup V\cup T}A_{ij}^{(\ell)}},$ (8)
$\alpha_{v\to t}^{(\ell)}$ $=\frac{\sum_{i\in V}\sum_{j\in T}A_{ij}^{(\ell)}}{\sum_{i\in V}\sum_{j\in S\cup V\cup T}A_{ij}^{(\ell)}},$
where $A_{ij}^{\ell}$ denotes the softmax-normalized attention weight from query token $i$ to key token $j$ at layer $\ell$; $S$, $V$, and $T$ represent the system prompt, visual, and textual tokens, respectively. To further enhance efficiency, all visual tokens are removed at a certain layer $\ell$ if and only if both $\alpha_{t\to v}^{(\ell)}$ and $\alpha_{v\to t}^{(\ell)}$ are lower than a threshold $\tau$. A more detailed ablation of the threshold can be found in Appendix. By removing all visual tokens at this point, the model further avoids redundant visual computation in the remaining prefilling and decoding stages, yielding slight improvements in both efficiency and performance.
@@ 5 Experiments
Method MME ScienceQA GQA POPE MMBench-EN MMBench-CN VizWiz VQAv2 Average
Upper Bound, 2880 Tokens
[…]
Ratio=25%, Retain up to 720 Tokens
FastV [chen2024fastv] 1477.3 69.8 60.4 83.1 65.6 55.4 57.2 77.2 95.4%
SparseVLM [zhang2024sparsevlm] 1446.1 67.5 60.9 71.0 63.8 55.4 58.6 77.2 93.1%
FasterVLM [zhang2024fastervlm] 1454.6 67.1 61.3 87.2 66.0 56.8 58.4 76.4 96.0%
GlobalCom2 [liu2025globalcom2] 1468.7 68.1 61.4 87.6 64.0 54.4 58.7 76.6 95.6%
DivPrune [alvar2025divprune] 1486.5 70.0 61.8 87.4 64.6 56.4 58.5 76.4 96.6%
ToDRE (Ours) 1504.3 70.7 63.3 87.5 66.6 58.0 59.5 77.5 98.2%
Ratio=10%, Retain up to 288 Tokens
FastV [chen2024fastv] 1282.9 69.3 55.9 71.7 61.6 53.5 56.1 70.2 88.8%
SparseVLM [zhang2024sparsevlm] 1332.2 68.6 56.1 63.2 54.5 52.3 56.2 69.9 86.3%
FasterVLM [zhang2024fastervlm] 1359.2 66.5 56.9 83.6 61.6 55.1 55.6 72.3 91.4%
GlobalCom2 [liu2025globalcom2] 1365.5 68.7 57.1 83.8 61.8 55.0 56.6 71.8 92.2%
DivPrune [alvar2025divprune] 1396.2 69.3 59.2 84.3 63.1 55.7 56.6 73.2 93.5%
ToDRE (Ours) 1464.4 70.2 59.4 85.0 63.3 57.8 56.7 74.4 95.0%
[…]
Ratio=25%, Retain up to 720 Tokens
FasterVLM [zhang2024fastervlm] 1516.1 71.1 62.3 86.1 67.6 62.1 58.1 76.1 95.6%
GlobalCom2 [liu2025globalcom2] 1531.2 71.4 62.7 86.5 67.9 61.3 58.2 77.2 96.0%
DivPrune [alvar2025divprune] 1530.2 71.4 62.9 87.0 67.7 61.4 60.6 77.4 96.6%
ToDRE (Ours) 1557.0 72.8 63.8 87.5 69.1 63.9 57.6 78.5 97.3%
Ratio=10%, Retain up to 288 Tokens
FasterVLM [zhang2024fastervlm] 1386.2 70.5 58.1 81.6 61.7 53.5 55.9 77.1 89.1%
GlobalCom2 [liu2025globalcom2] 1399.5 71.0 58.3 82.4 65.0 56.6 55.6 72.8 90.8%
DivPrune [alvar2025divprune] 1463.3 70.7 60.1 86.5 64.3 53.0 59.1 75.4 92.5%
ToDRE (Ours) 1490.2 71.4 59.9 83.7 65.3 60.5 56.9 75.9 93.6%
Table 1: Performance of training-free token compression methods across eight image-language benchmarks. “Average” denotes the mean performance ratio between each token compression method and the vanilla LLaVA-NeXT-7B. We evaluate all methods at retention ratios of 25% and 10%, with the best results highlighted in bold.
[…]
FastV [chen2024fastv] 25% 720 32.3 31.2 16.5 40.3 90.4%
FasterVLM [zhang2024fastervlm] 33.8 41.0 19.6 40.5 102.3%
DivPrune [alvar2025divprune] 33.6 41.8 19.5 40.4 102.5%
ToDRE (Ours) 33.3 42.4 19.6 41.0 103.1%
FastV [chen2024fastv] 10% 288 30.4 30.4 11.2 38.7 80.8%
FasterVLM [zhang2024fastervlm] 34.3 36.1 19.1 36.9 96.4%
DivPrune [alvar2025divprune] 33.3 40.9 18.9 40.3 100.7%
ToDRE (Ours) 33.2 41.9 18.2 40.7 100.9%
Qwen2.5-VL-7B-Instruct [bai2025qwen25vl] InternVL2-8B [internvl2]
Benchmark Original Ret. 25% Ret. 10% Original Ret. 25% Ret. 10%
MME [Fu2023mme] 1687.7 1680.6 1573.9 1628.8 1566.4 1424.4
ScienceQA [lu2022sqa] 88.5 86.4 85.4 96.5 95.3 92.5
GQA [hudson2019gqa] 60.9 57.3 52.9 62.8 56.9 52.1
POPE [li2023pope] 87.7 85.4 80.8 87.8 83.7 76.9
MMBench-EN [Liu2023mmbench] 82.9 79.9 72.8 81.2 77.2 71.4
MMBench-CN [Liu2023mmbench] 81.7 78.0 70.4 80.0 74.2 67.6
VizWiz [gurari2018vizwiz] 70.6 68.7 66.7 60.6 58.6 56.3
VQAv2 [goyal2017vqav2] 82.9 78.3 72.8 79.0 72.8 68.6
VideoMME [fu2024videomme] 61.5 59.4 57.3 55.0 54.2 52.6
Egoschema [mangalam2023egoschema] 58.3 57.3 55.8 55.9 55.0 52.5
MLVU [zhou2024mlvu] 59.3 58.7 57.2 47.3 52.8 54.0
LongVideoBench [wu2024longvideobench] 58.4 57.7 54.8 55.0 52.3 48.8
Average 100.0% 97.1% 92.0% 100.0% 96.8% 91.5%
Table 3: Performance of ToDRE on Qwen2.5-VL-7B-Instruct and InternVL2-8B. Benchmarks as rows. “Ret.”=Retention Ratio. Averages are normalized to each model’s Original (=100%).
@@
We evaluate ToDRE over multiple prevalent LVLMs (including LLaVA-NeXT-7B/13B [liu2024llavanext], Qwen2.5-VL-7B-Instruct [bai2025qwen25vl], and InternVL2-8B [internvl2]) and twelve widely adopted benchmarks (including eight on image understanding tasks and four on video understanding tasks). More details on the benchmarks, network backbones, and comparison methods can be found in the Appendix.
5.1 Benchmarking
@@
In Table˜1, we report ToDRE’s performance on a range of image‐understanding benchmarks at different token-retention ratios. First, under the same setup where 75% of visual tokens are pruned in Stage 1—matching competing methods—ToDRE further removes all remaining visual tokens in Stage 2 and achieves a 98.2% average score, outperforming the second-best method by 1.6%. Second, under more extreme compression (only 10% of visual tokens are retained), ToDRE surpasses the second-best approach by 1.5%. Third, ToDRE also achieves top performance on larger models, reaching an average score of 93.6% on the 13B variant—demonstrating strong adaptability across model scales. Note that FastV [chen2024fastv] and SparseVLM [zhang2024sparsevlm] are excluded from the 13B comparison, as their pruning strategies, originally tailored for the 7B model, lead to substantial performance degradation when directly transferred to the 13B model. This further underscores the robustness and transferability of ToDRE.
@@
To further assess ToDRE’s generalization ability, we evaluate it on both short- and long-form video understanding benchmarks. As shown in Table˜2, ToDRE outperforms the baseline by 3.1% and 0.9% under the same token retention ratios used for images, and surpasses the second-best method by 0.6% and 0.2%, respectively. Interestingly, ToDRE even surpasses the baseline model in some cases. We attribute this to the reduced interference from redundant visual tokens, which may otherwise suppress task-relevant information during inference. Similarly, SparseVLM is excluded due to transferability issues, and GlobalCom2 [liu2025globalcom2] is omitted as it is specifically designed for image-only inputs. In contrast, ToDRE demonstrates broad generalization across both modalities and model scales.
Cross-Model Evaluation.
As shown in Table˜3, we further evaluate ToDRE on Qwen and InternVL backbones. Specifically, ToDRE retains 97.1% and 96.8% of the original performance on Qwen2.5-VL-7B-Instruct and InternVL2-8B at a 25% retention ratio, respectively, and still maintains more than 90% of the original performance even when only 10% of visual tokens are preserved, demonstrating strong robustness across different model architectures.
5.2 Efficiency
[…]